The sum of the ages of Armando and Shantel is 72 years. 9 years ago, Armando's age was 2 times Shantel's age. How old is Armando now?
Answer by addingup(3677)

Armando 9 years ago was 2 x (Shantel - 9)
A + S = 72
subtract S from both sides:
A = 72 - S
-------------------------------
A - 9 = 2(S-9)
A - 9 = 2S - 18 substitute for A:
72 - S - 9 = 2S - 18
63 - S = 2S - 18
81 = 3S flip it so we can get the unknown on the left:
3S = 81
S = 81/3
S = 27
A = 72-27 = 45